# Gravitational Atoms from Topological Stars

### v1

@@ III.4 Charged modes and gravitating Thomson dipoles
The results are displayed in Fig. 14, which we discuss next. As a first observation, we see that in both neutral and charged configurations (left and right columns, respectively) the three types of modes introduced in Section III.2 are clearly distinguished: TS modes (top row) are localized near the surface of the bubble, mixed modes (middle row) probe the surface of the bubble but are localized at a larger radius, and cloud modes (bottom row) localize entirely in a region that does not intersect the bubble. As advanced earlier, the charged fields for our choice of parameters (in particular $m_{\text{c}}=N_{\text{c}}/2$) correspond to north modes, which exist at the northern $z$-axis but vanish at the southern one (a south mode, corresponding to $m_{\text{c}}=-N_{\text{c}}/2$, would have the same form but upside-down). Given that they are localized in a smaller region of space than the neutral modes but carry the exact same amount of energy by construction, the maximum field amplitude is (roughly ten times) higher in the charged case. Regarding the angular momentum, in agreement with (77) we see that the charged modes carry a larger amount. In the neutral case, the angular momentum is of orbital nature, i.e. due to the rotation of the field around the bubble. The situation is less intuitive for charged fields. Heuristically, the origin of the angular momentum in that case can be explained from the properties of the Thomson dipole Thomson (2009). The latter is simply a pair of point charges, one electric and one magnetic, which are separated and at rest. Such a static configuration has a non-trivial pointing vector, which endows the electromagnetic field with an angular momentum proportional to the product of the charges, and pointing from the electric to the magnetic charge (see Garfinkle and Rey (1991); Bunster and Henneaux (2007); Dyson and Pereñiguez (2023); Pereñiguez et al. (2024) for the BH generalization). With this picture in mind, it is clear that the charged field configurations in the second column of Fig. 14 are simply a wave-like version of a Thomson dipole, where the bubble plays the role of the magnetic charge and the field replaces the electric charge. Consequently, these states carry angular momentum along the $z$-direction, even though they localize along the $z$-axis.
@@
Figure 14: Field amplitudes $\lvert\varphi_{\text{n,c}}\rvert$ in the plane spanned by $\tilde{x}\equiv\tilde{r}\sin\theta$ (horizontal axis) and $\tilde{z}\equiv\tilde{r}\cos\theta$ (vertical axis), for a TS background with $\eta=1.75$ and a scalar field mass $\tilde{\mu}=7$ (matching Fig. 6). The bubble’s surface (dashed black circle) is at $\tilde{r}=\sqrt{\tilde{x}^{2}+\tilde{y}^{2}}=1$, and the coordinate ranges are adapted to capture the cloud’s effective size. As explained in the main text, the wavefunctions are normalized by fixing their total energies to $E[\varphi_{\text{n,c}}]/r_{B}=10^{-3}$, and constant-amplitude contours (solid black lines) are included to guide the eye. Each panel shows the fundamental mode with the indicated values of $N_{\text{n,c}},\ell_{\text{n,c}},m_{\text{n,c}}$, and displays the corresponding values of $\Lambda_{\text{n,c}}$ and total (dimensionless) angular momentum $L_{\text{n,c}}\equiv L[\varphi_{\text{n,c}}]/r_{B}^{2}$. Left and right columns correspond to neutral ($N_{\text{n}}=0$) and charged configurations, respectively. Rows share identical frequencies (which can be obtained from Fig. 6) and radial profiles, and correspond to TS modes (a)-(b), mixed modes (c)-(d) and cloud modes (e)-(f).
@@ IV Discussion
We have studied the structure of massive bound states in the vicinity of TSs, focusing on charged scalars minimally coupled to the Einstein–Maxwell theory. First, we have established generic properties expected to hold for other compact objects. Namely, that bound states are strictly normal, and the regimes $r_{b}/\lambda\ll 1$ and $r_{b}/\lambda\gg 1$ are hydrogen-like and governed by congruences of timelike geodesics, respectively. Next, we have proposed a bound-mode classification based on the classically allowed regions that modes enjoy, and their position relative to the ISCO. In the case of TSs, considering modes with no KK momentum, modes fall in three categories: TS-modes that exist close to the star’s center, cloud modes that exist beyond the ISCO, and mixed modes that experience two classically allowed regions separated by a centrifugal barrier. If KK momentum is included, additional potential barriers appear and the mode classification becomes richer. In particular, we have found that modes with too large KK momentum cannot remain bound to the star, and can only be radiative. In the case that the field is electrically charged, we have shown that, while the spectrum and radial distribution resembles closely (or is even identical to) the neutral case, their spatial configuration changes drastically. In fact, we have shown that TSs and charged-field bound states provide a smooth horizonless realization of a gravitating Thomson dipole, so far only constructed with BHs.
@@
This work shows that TSs yield genuine realizations of a gravitational atom, superseding BH ones in that they are truly stationary, linearly-stable systems. However, unlike for BHs where clouds grow spontaneously due to superradiant instabilities, it is unclear how TS gravitational atoms may form. One possibility consists in assuming a non-zero scalar-field background density, motivated from dark matter considerations Hui et al. (2019); Clough et al. (2019), so that the massive states described above would unavoidably form around TSs.
@@
A mechanism with less assumptions would be that TS gravitational atoms form spontaneously, and correspond to end states of instabilities, in closer analogy to the BH case. However, it is unlikely that the instability is of superradiant nature – a fact that is in agreement with the intuition that TSs correspond to microscopic fundamental states. The reason is that in the absence of an horizon there is no dissipation channel in the scattering of waves, which is one of the fundamental ingredients of superradiance Zel’Dovich (1972); Bekenstein and Schiffer (1998). This is for instance the case of ideal fluid stars, which do not experience superradiance and dissipative effects such as conductivity Richartz and Saa (2013); Cardoso et al. (2017) or viscosity Redondo-Yuste and Cardoso (2025) need to be accounted for in order to obtain it. The absence of superradiance in the spherical TSs considered here follows easily: massless modes at infinity behave as $\psi(r)\sim e^{-i\omega r}+Re^{i\omega r}$, corresponding to incoming and outgoing waves, where $R$ is the reflectivity coefficient. If the frequency is real, then the regularity condition (17) implies that $\psi(r)$ and $\bar{\psi}(r)$ are linearly dependent solutions, and by the vanishing of the Wronskian one finds $\lvert R\rvert=1$, corresponding to perfect reflectivity so, in particular, there is no amplification. A similar argument is expected to hold for rotating TSs, only recently constructed in Bianchi et al. (2025a); Heidmann et al. (2025), as long as a regularity condition similar to (17) holds at the star’s center. However, it would be interesting to explore whether the electromagnetic coupling can change this picture, as it has been shown to yield new regions of negative-energy states and different phenomenology in the BH case Dyson and Pereñiguez (2023); Pereñiguez et al. (2024).
@@
Alternatively, in Heidmann et al. (2025) the authors notice that the rotating solution presents a 5-dimensional ergoregion. This indicates that KK-modes might be subject to a different class of instability not based on superradiant scattering, known as ergoregion instability Friedman (1978) (see e.g. Cardoso et al. (2006) in the context of microstate geometries). This is associated to spaces with ergoregions, but no horizons, where a field grows unbounded as long as it is in a negative energy state and it radiates to infinity. We have shown that massive modes with large-enough KK momentum are necessarily radiative. Assuming that this is also true for the rotating TS, combined with the 5-dimensional ergoregion of Heidmann et al. (2025), it follows that massive modes with enough KK momentum can be radiative and simultaneously occupy negative-energy states, thus meeting the assumptions that lead to an ergoregion instability. Understanding whether this instability can yield, as an end state, a TS with a stationary bound-state cloud such as those considered here requires tracking non-linearly the evolution of the instability. The knowledge of explicit rotating TS solutions Bianchi et al. (2025a); Heidmann et al. (2025) allows a precise analysis of these questions, at least at the perturbative level, where one can sharply test the expectations discussed here.
@@ Appendix A Wu-Yang monopole harmonics
The Wu-Yang monopole harmonics $Y_{N,\ell,m}$ Wu and Yang (1976) are eigen-functions of a well-defined set of angular-momentum operators, in the background of a magnetic monopole. They live in a $U(1)$-line bundle, in the sense that under a gauge-transformation (6) they transform as $Y_{N,\ell,m}\mapsto e^{i\alpha}Y_{N,\ell,m}$. This point is reviewed explicitly in Pereñiguez et al. (2024), and here we simply write the resulting explicit expressions for $Y_{N,\ell,m}$:
@@
It is useful to note the characteristic problem $Y_{N,\ell,m}$ are solutions to. Consider the Euclidean 3-dimensional space, a monopole field $A=-P\cos\theta d\phi$, and introduce the gauge-covariant operators,
@@ Appendix B Equations of Motion and Bianchi Identities
and one can show that the following identities hold,
@@ Appendix C Quasimodes in the eikonal limit
Quasimodes are approximate solutions to the wave equation localized at the minimum of the potential. They allow us to determine the real part of the frequency in an appropriate limit. The way we construct them here, they are insensitive to the imaginary part of the frequency. That vanishes for bounded modes on a TS background as shown in the main text. However, the same method can also be applied to the black string regime, which we do not show here, but we have checked that the method works well. In order to simplify the algebra, we use the dimensionless coordinates, (42), together with the following radial variable:
@@
First, there is a local minimum near $z_{0}=0$, but we need to determine higher order terms in its eikonal expansion in order to talk about its position relative to the star surface. To that end, we need to look for the minimum simultaneously with solving the radial equation so as to obtain higher order $w_{j}$ pieces. That is rather tedious and we just state the result - namely, this extremum does not sit outside of the star surface/event horizon when corrections are included. Thus, we will not focus on it here.
@@
Finally, we have a local minimum for the outermost extremum - the one we are interested in. The choice of prefactor in (89) was made intentionally, so that our quasimode construction is in the vicinity of $z=1+\mathcal{O}(\ell^{-1})$ for $\ell\gg 1$.
@@
To solve (90) in the eikonal regime, we take:
@@
where $c_{\phi}$ is a constant related to the amplitude of the radial function. We have made a choice of sign [^9] It is here, where one can determine the sign of $w_{2}$, before the location of the potential minimum we are interested in has been clarified. It is only after that we can decide which $\phi(z)$ solution to pick. for $\phi(z)$, so that $e^{-\ell\,\phi(z)}$ decays away from $z=1$. From here onward, we can keep integrating the resulting equation at every order, looking for a smooth solution, in order to determine $w_{j}$ and $Z_{j}$. In particular, we get
@@
Inverting (89), we can immediately see that the location of the minimum, where the clouds get localized, and where we expect their peak to be, scales as $\tilde{r}\sim\ell^{2}$ for $\ell\gg 1$. We can also get an estimate for a cloud’s size in the eikonal regime by looking at the second derivative of the approximate radial wavefunction, (93), in terms of $\tilde{r}$. The outcome is:
@@ Appendix D Quasimodes in the large scalar mass limit
The computation in this section is valid for the TS case only, as it targets the modes that localize in the classically allowed region that extends from the star surface outwards in the language of Section (III.2) - that is, the TS modes.
One should be able to replicate the construction from the previous section, but targeting the TS modes, as qualitatively they are also bound states sharply peaked in a local minimum. From the regime of validity of the eikonal approximation [^10] All arguments in both approximations hold for $\tilde{k}=0$, where it is much easier to infer the relevant regime., one anticipates that $\tilde{\mu}$ much larger than any other scale in the problem should be the appropriate regime to characterize these modes, most likely in a regime where they are very close to the star surface.
The mathematical procedure is the same as in the previous section, just in a different limit, namely, $\tilde{\mu}\gg 1$. We first look at the derivative of the potential in order to identify a change of coordinates, so that the minimum is localized at $z=1$ to leading order for $\tilde{\mu}\gg 1$. The required transformation is:
@@
where as in the eikonal approximation, $\hat{Z}_{0}$ is just a constant that can be reabsorbed in $\hat{\phi}(z)$, and $\tilde{\omega}_{0}$ is what we get from the determination of the leading order term in the location of the minimum near the star surface. The end result is:
@@ Appendix E Leaver’s method
In the large $n$ limit, the ratio of series coefficients in Eq. (106) behaves as $a_{n+1}/a_{n}\sim 1\pm\tfrac{\sqrt{2(\eta-1)\gamma/\eta}}{\sqrt{n}}+\mathcal{O}(n^{-1})$. Recalling that $\gamma>0$ for bound modes, the series converges uniformly if we pick the minus branch of the solution. This branch corresponds to minimal solutions of the radial equation Leaver (1985), with the ratio of successive $a_{n}$ coefficients given by the infinite continued fraction
@@
$\left(\beta_{n}-\frac{\alpha_{n-1}\delta_{n}}{\beta_{n-1}-}...\frac{\alpha_{0}\delta_{1}}{\beta_{1}}\right)=\frac{\alpha_{n}\delta_{n+1}}{\beta_{n+1}-}\frac{\alpha_{n+1}\delta_{n+2}}{\beta_{n+2}-}$. (110)

### v2

@@ III.4 Charged modes and gravitating Thomson dipoles
The results are displayed in Fig. 14, which we discuss next. As a first observation, we see that in both neutral and charged configurations (left and right columns, respectively) the three types of modes introduced in Section III.2 are clearly distinguished: TS modes (top row) are localized near the surface of the bubble, mixed modes (middle row) probe the surface of the bubble but are localized at a larger radius, and cloud modes (bottom row) are localized entirely in a region that does not intersect the bubble. As anticipated earlier, the charged fields for our choice of parameters (in particular $m_{\text{c}}=N_{\text{c}}/2$) correspond to north modes, which exist at the northern $z$-axis but vanish at the southern one (a south mode, corresponding to $m_{\text{c}}=-N_{\text{c}}/2$, would have the same form but upside-down). Given that they are localized in a smaller region of space than the neutral modes but carry the exact same amount of energy by construction, the maximum field amplitude is (roughly ten times) higher in the charged case. Regarding the angular momentum, in agreement with Eq. (77) we see that the charged modes carry a larger amount. In the neutral case, the angular momentum is of orbital nature, i.e. due to the rotation of the field around the bubble. The situation is less intuitive for charged fields. Heuristically, the origin of the angular momentum in that case can be explained from the properties of the Thomson dipole Thomson (2009). The latter is simply a pair of point charges, one electric and one magnetic, which are separated and at rest. Such a static configuration has a non-trivial Poynting vector, which endows the electromagnetic field with an angular momentum proportional to the product of the charges, and pointing from the electric to the magnetic charge (see Garfinkle and Rey (1991); Bunster and Henneaux (2007); Dyson and Pereñiguez (2023); Pereñiguez et al. (2024) for the BH generalization). With this picture in mind, it is clear that the charged field configurations in the second column of Fig. 14 are simply a wave-like version of a Thomson dipole, where the bubble plays the role of the magnetic charge and the field replaces the electric charge. Consequently, these states carry angular momentum along the $z$-direction, even though they are localized along the $z$-axis.
@@
Figure 14: Field amplitudes $\lvert\varphi_{\text{n,c}}\rvert$ in the plane spanned by $\tilde{x}\equiv\tilde{r}\sin\theta$ (horizontal axis) and $\tilde{z}\equiv\tilde{r}\cos\theta$ (vertical axis), for a TS background with $\eta=1.75$ and a scalar field mass $\tilde{\mu}=7$ (matching Fig. 6). The bubble’s surface (dashed black circle) is at $\tilde{r}=\sqrt{\tilde{x}^{2}+\tilde{y}^{2}}=1$, and the coordinate ranges are adapted to capture the cloud’s effective size. As explained in the main text, the wavefunctions are normalized by fixing their total energies to $E[\varphi_{\text{n,c}}]/r_{B}=10^{-3}$, and constant-amplitude contours (solid black lines) are included to guide the eye. Each panel shows the fundamental mode with the indicated values of $N_{\text{n,c}},\ell_{\text{n,c}},m_{\text{n,c}}$, and displays the corresponding values of $\Lambda_{\text{n,c}}$ and total (dimensionless) angular momentum $L_{\text{n,c}}\equiv L[\varphi_{\text{n,c}}]/r_{B}^{2}$. Left and right columns correspond to neutral ($N_{\text{n}}=0$) and charged configurations, respectively. Rows share identical frequencies (which can be obtained from Fig. 6) and radial profiles, and correspond to TS modes (panels a, b), mixed modes (panels c, d) and cloud modes (panels e, f).
@@ IV Discussion
We have studied the structure of massive bound states in the vicinity of TSs, focusing on charged scalars minimally coupled to the Einstein–Maxwell theory. First, we have established generic properties expected to hold for other compact objects – namely, that bound states are strictly normal, and the regimes $r_{b}/\lambda\ll 1$ and $r_{b}/\lambda\gg 1$ are hydrogen-like and governed by congruences of timelike geodesics, respectively. Next, we have proposed a bound-mode classification based on the classically allowed regions for the modes, and their position relative to the ISCO. In the case of TSs, considering modes with no KK momentum, modes fall in three categories: TS modes that exist close to the star’s center, cloud modes that exist beyond the ISCO, and mixed modes that experience two classically allowed regions separated by a centrifugal barrier. If KK momentum is included, additional potential barriers appear and the mode classification becomes richer. In particular, we have found that modes with too large KK momentum cannot remain bound to the star, and can only be radiative. In the case that the field is electrically charged, we have shown that, while the spectrum and radial distribution resemble closely (or are even identical to) the neutral case, their spatial configuration changes drastically. In fact, we have shown that TSs and charged-field bound states provide a smooth horizonless realization of a gravitating Thomson dipole, so far only constructed with BHs.
@@
This work shows that TSs yield genuine realizations of a gravitational atom, superseding BH ones in that they are truly stationary, linearly stable systems. However, unlike for BHs where clouds grow spontaneously due to superradiant instabilities, it is unclear how TS gravitational atoms may form. One possibility consists in assuming a non-zero scalar-field background density, motivated from dark matter considerations Hui et al. (2019); Clough et al. (2019), so that the massive states described above would unavoidably form around TSs.
@@
A mechanism with less assumptions would be that TS gravitational atoms form spontaneously, and correspond to end states of instabilities, in closer analogy to the BH case. However, it is unlikely that the instability is of superradiant nature – a fact that is in agreement with the intuition that TSs correspond to microscopic fundamental states. The reason is that in the absence of an horizon there is no dissipation channel in the scattering of waves, which is one of the fundamental ingredients of superradiance Zel’Dovich (1972); Bekenstein and Schiffer (1998). This is for instance the case of ideal fluid stars, which do not experience superradiance, and dissipative effects such as conductivity Richartz and Saa (2013); Cardoso et al. (2017) or viscosity Redondo-Yuste and Cardoso (2025) need to be accounted for in order to trigger superradiance. The absence of superradiance in the spherical TSs considered here follows easily: massless modes at infinity behave as $\psi(r)\sim e^{-i\omega r}+Re^{i\omega r}$, corresponding to incoming and outgoing waves, where $R$ is the reflectivity coefficient. If the frequency is real, then the regularity condition (17) implies that $\psi(r)$ and $\bar{\psi}(r)$ are linearly dependent solutions, and by the vanishing of the Wronskian one finds $\lvert R\rvert=1$, corresponding to perfect reflectivity – so, in particular, there is no amplification. A similar argument is expected to hold for rotating TSs, only recently constructed in Bianchi et al. (2025a); Heidmann et al. (2025), as long as a regularity condition similar to Eq. (17) holds at the star’s center. However, it would be interesting to explore whether the electromagnetic coupling can change this picture, as it has been shown to yield new regions of negative-energy states and different phenomenology in the BH case Dyson and Pereñiguez (2023); Pereñiguez et al. (2024).
@@
Alternatively, in Ref. Heidmann et al. (2025) the authors notice that the rotating solution presents a 5-dimensional ergoregion. This indicates that KK-modes might be subject to a different class of instability not based on superradiant scattering, known as ergoregion instability Friedman (1978) (see e.g. Cardoso et al. (2006) in the context of microstate geometries). This is associated to spaces with ergoregions, but no horizons, where a field grows unbounded as long as it is in a negative energy state and it radiates to infinity. We have shown that massive modes with large-enough KK momentum are necessarily radiative. Assuming that this is also true for the rotating TS, combined with the 5-dimensional ergoregion of Ref. Heidmann et al. (2025), it follows that massive modes with enough KK momentum can be radiative and simultaneously occupy negative-energy states, thus meeting the assumptions that lead to an ergoregion instability. Understanding whether this instability can yield, as an end state, a TS with a stationary bound-state cloud such as those considered here requires tracking non-linearly the evolution of the instability. The knowledge of explicit rotating TS solutions Bianchi et al. (2025a); Heidmann et al. (2025) allows a precise analysis of these questions, at least at the perturbative level, where one can sharply test the expectations discussed here.
@@ Appendix A Wu-Yang monopole harmonics
The Wu-Yang monopole harmonics $Y_{N,\ell,m}$ Wu and Yang (1976) are eigenfunctions of a well-defined set of angular momentum operators, in the background of a magnetic monopole. They live in a $U(1)$-line bundle, in the sense that under a gauge-transformation (6) they transform as $Y_{N,\ell,m}\mapsto e^{i\alpha}Y_{N,\ell,m}$. This point is reviewed explicitly in Pereñiguez et al. (2024), and here we simply write the resulting explicit expressions for $Y_{N,\ell,m}$:
@@
It is useful to note the characteristic problem that the $Y_{N,\ell,m}$’s solve. Consider the Euclidean 3-dimensional space, a monopole field $A=-P\cos\theta d\phi$, and introduce the gauge-covariant operators,
@@ Appendix B Equations of Motion and Bianchi Identities
and one can show that the following identities hold:
@@ Appendix C Quasimodes in the eikonal limit
Quasimodes are approximate solutions to the wave equation localized at the minimum of the potential. They allow us to determine the real part of the frequency in an appropriate limit. As we construct them here, they are insensitive to the imaginary part of the frequency, that vanishes for bounded modes on a TS background (as shown in the main text). However, the same method can also be applied to the black string regime, which we do not show here, but we have checked that the method works well. In order to simplify the algebra, we use the dimensionless coordinates of Eq. (42), together with the following radial variable:
@@
First, there is a local minimum near $z_{0}=0$, but we need to determine higher order terms in its eikonal expansion in order to talk about its position relative to the star surface. To that end, we need to look for the minimum while simultaneously solving the radial equation so as to obtain higher order $w_{j}$ pieces. That is rather tedious and we just state the result – namely, this extremum does not sit outside of the star surface/event horizon when corrections are included. Thus, we will not focus on it here.
@@
Finally, we have a local minimum for the outermost extremum. This is the one we are interested in. The choice of prefactor in (89) was made intentionally, so that our quasimode construction is in the vicinity of $z=1+\mathcal{O}(\ell^{-1})$ for $\ell\gg 1$.
@@
To solve Eq. (90) in the eikonal regime, we take:
@@
where $c_{\phi}$ is a constant related to the amplitude of the radial function. We have made a choice of sign for $\phi(z)$, so that $e^{-\ell\,\phi(z)}$ decays away from $z=1$. (It is here that one can determine the sign of $w_{2}$, before the location of the potential minimum we are interested in has been clarified. It is only after that we can decide which $\phi(z)$ solution to pick.) From here onward, we can keep integrating the resulting equation at every order, looking for a smooth solution, in order to determine $w_{j}$ and $Z_{j}$. In particular, we get
@@
Inverting Eq. (89), we can immediately see that the location of the minimum, where the clouds get localized, and where we expect their peak to be, scales as $\tilde{r}\sim\ell^{2}$ for $\ell\gg 1$. We can also get an estimate for the cloud’s size in the eikonal regime by looking at the second derivative of the approximate radial wavefunction (93) in terms of $\tilde{r}$. The outcome is:
@@ Appendix D Quasimodes in the large scalar mass limit
The computation in this section is valid for the TS case only, as it targets modes localized in the classically allowed region that extends from the star surface outwards in the language of Section III.2 – that is, the TS modes.
One should be able to replicate the construction from the previous section, but targeting the TS modes, as qualitatively they are also bound states sharply peaked in a local minimum. From the regime of validity of the eikonal approximation (all arguments in both approximations hold for $\tilde{k}=0$, where it is much easier to infer the relevant regime), one anticipates that the appropriate regime to characterize these modes is one where $\tilde{\mu}$ is much larger than any other scale in the problem – most likely, a regime where the modes are very close to the star’s surface.
The mathematical procedure is the same as in the previous section, just in a different limit, namely, $\tilde{\mu}\gg 1$. We first look at the derivative of the potential in order to identify a change of coordinates, so that the minimum is localized at $z=1$ to leading order in $\tilde{\mu}\gg 1$. The required transformation is:
@@
where as in the eikonal approximation, $\hat{Z}_{0}$ is just a constant that can be reabsorbed in $\hat{\phi}(z)$, and $\tilde{\omega}_{0}$ is what we get from the determination of the leading-order term in the location of the minimum near the star surface. The end result is:
@@ Appendix E Leaver’s method
In the large-$n$ limit, the ratio of series coefficients in Eq. (106) behaves as $a_{n+1}/a_{n}\sim 1\pm\tfrac{\sqrt{2(\eta-1)\gamma/\eta}}{\sqrt{n}}+\mathcal{O}(n^{-1})$. Recalling that $\gamma>0$ for bound modes, the series converges uniformly if we pick the minus branch of the solution. This branch corresponds to minimal solutions of the radial equation Leaver (1985), with the ratio of successive $a_{n}$ coefficients given by the infinite continued fraction
@@
$\left(\beta_{n}-\frac{\alpha_{n-1}\delta_{n}}{\beta_{n-1}-}...\frac{\alpha_{0}\delta_{1}}{\beta_{1}}\right)=\frac{\alpha_{n}\delta_{n+1}}{\beta_{n+1}-}\frac{\alpha_{n+1}\delta_{n+2}}{\beta_{n+2}-}$.\,, (110)